OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!!!!!!!!! a^4+4b^4

OpenStudy (mathmale):

Instructions, please?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

thats all it says ;p

OpenStudy (mathmale):

Suggestions: Let A=a^2 and B=b^2. Please make these substitution now. Are you certain that that sign in the middle is + and not -?

OpenStudy (anonymous):

you cannot factor the sum of two fourth powers

OpenStudy (anonymous):

It's a positive

OpenStudy (vishweshshrimali5):

Actually I think you may...

OpenStudy (anonymous):

at least not with real numbers

OpenStudy (vishweshshrimali5):

Do one thing.. add and subtract 4*a^2 * b^2 and then group +4a^2 b^2 with the a^4 + 4b^4

OpenStudy (anonymous):

oh am i ever WRONG sorry

OpenStudy (mathmale):

Then, unfortunately, you cannot find real factors; you could factor this, but your results will be complex (part imaginary). Are you familiar with complex numbers? imaginary numbers?

OpenStudy (anonymous):

do what @vishweshshrimali5 is suggesting, it will work

OpenStudy (vishweshshrimali5):

Okay... let me give the solution.. part of it... you can work out the rest

OpenStudy (vishweshshrimali5):

\[a^4 + 4b^4\] \[=a^4 + 4b^4 + 4a^2 b^2 - 4a^2 b^2\] \[=(a^2)^2 + (2b^2)^2 + 2*(a^2)*(2b^2) - 4a^2b^2\] \[=(a^2)^2 + (2b^2)^2 + 2*(a^2)*(2b^2) - (2ab)^2\] \[=((a^2)^2 + (2b^2)^2 + 2*(a^2)*(2b^2)) - (2ab)^2\]

OpenStudy (vishweshshrimali5):

Now, first use this formula: \[x^2 +2xy + y^2 = (x+y)^2\] for the first bracket thing: \(a^2 + (2b)^2 + 2(a^2)(2b^2)\) What will you get?

OpenStudy (anonymous):

think this method of completing the square is called the "sophie germain" trick

OpenStudy (vishweshshrimali5):

Wow! Never knew that :D

OpenStudy (anonymous):

she pretended to be a man during the french revolution, got instructions from legrange

OpenStudy (vishweshshrimali5):

Now, if only I knew that story -.- Maths is definitely interesting.. and I see proofs of that every single day :D

OpenStudy (mathmale):

Let's not forget that this is Maddiem's question, and she deserves to be involved in this discussion.

OpenStudy (anonymous):

i friend of mine wrote some notes for a class, included that as information i assume he was right, did not check, but i imagine a quick google search would get it

OpenStudy (vishweshshrimali5):

It did ;)

OpenStudy (anonymous):

a google search actually calls this question "germain's identity"!!

OpenStudy (yttrium):

yes ^_^ and the identity is exactly this given question!

OpenStudy (mathmale):

Please note: Maddiem, the person who posted this question, has left (gone offline). I admire your ability to do this particular math (I would not have thought about such an approach), but urge you not to forget that a student was waiting for your attention and help.

OpenStudy (anonymous):

Sorry everyone for leaving my computer died ;P Thank you so much for all the help :*. I'll fan all of you!!

OpenStudy (anonymous):

you got this?